OpenStudy (promisespromises):

find the point on the graph of y=1+1/x at which the tangent line passes through the point (3,0)

OpenStudy (anonymous):

Let \((x_0,y_0)\) be the intersection of the given curve and the tangent line at \(x=x_0\). The tangent line equation is \[y-y_0=y'(x_0)(x-x_0)\\ y-\left(1+\frac{1}{x_0}\right)=-\frac{1}{(x_0)^2}(x-x_0)\\ \color{red}{y=-\frac{1}{(x_0)^2}(x-x_0)+\left(1+\frac{1}{x_0}\right)}.\] It passes through the point (3, 0), so you know it can also be written as \[y-0=m(x-3)\\ y=m(x-3).\] Since these two equations represent the same line, they must have the same slope, so the last equation can be written as \[\color{blue}{y=-\frac{1}{(x_0)^2}(x-3)}.\] Set the red equation equal to the blue, and solve for \(x_0\). This gives you the x-coordinate of the point you're looking for, so you'll also have to solve for \(y_0\) by finding \(1+\frac{1}{x_0}\).

OpenStudy (promisespromises):

Wow thanks so much. I was completely stumped on this one.

OpenStudy (anonymous):

Actually, I may have made some computational mistakes at some point (in my own work on paper)... I'll go over this again. But the procedure is right.

OpenStudy (promisespromises):

Alright, thanks again.

OpenStudy (anonymous):

You're welcome

OpenStudy (promisespromises):

For some reason I can't get it to turn out right when I set them equal to each other. I keep getting weird values. Should I get an exact number or should it still have the x variable in it?

OpenStudy (anonymous):

The x should cancel. \[\color{red}{-\frac{1}{(x_0)^2}(x-x_0)+\left(1+\frac{1}{x_0}\right)}=\color{blue}{-\frac{1}{(x_0)^2}(x-3)}\] Multiplying both sides by \(-(x_0)^2\), you have \[x-x_0-(x_0)^2-x_0=x-3\\ -(x_0)^2-2x_0=-3\\ (x_0)^2+2x_0-3=0\]

OpenStudy (promisespromises):

Okay, thanks. I was making it harder than it had to be